OpenStudy (konradzuse):

find the third taylor polynomial for f(x) = 7/x expanded about c = 1

OpenStudy (konradzuse):

?

OpenStudy (anonymous):

we need to calculate \( f^n(a) \)

OpenStudy (anonymous):

a=1

OpenStudy (konradzuse):

it wouldn't be 7/x?

OpenStudy (anonymous):

\( f(1)=7 \) \( f'(1)=-7/1^2 =-7\) \( f''(1)=(-7)(-2)/1^3 =7 *2\) \( f'''(1)=(-7)(-2)(-3)/1^4 =-7 *2*3\) \(f^{(4)}(1)= (-7)(-2)(-3)(-4)/1^5=7*2*3*4\) ..... \(f^{(n)}(1)= (-7)(-2)(-3)(-4)...(-n)/1^{n+1}=(-1)^{n}*7*n!\)

OpenStudy (anonymous):